joe jogs 2.5 slower than kelly jogs. in the time that it takes joe to jog 30 miles, kathy can jog 40 miles. find the jogging speed of each person.
Answer by Alan3354(69443)

joe jogs 2.5 slower than kelly jogs. in the time that it takes joe to jog 30 miles, kathy can jog 40 miles. find the jogging speed of each person.
--------------
2.5 what?
If it's 2.5 mi/hr,
r = Joe's rate
-----
30/r = 40/(r+2.5)
30r + 75 = 40r
r = 7.5 mi/hr
Kelly's rate = 10 mi/hr